# Guido Franchetti1 and Derek Harland2

### v1

@@ 4 Killing spinors and pluricomplex structures
We extend complex conjugation to $\mathfrak{su}(2)$-valued spinors trivially, that is, for $s\in S$, $\upsilon\in\mathfrak{su}(2)$ we define $\overline{s\otimes\upsilon}=\overline{s}\otimes\upsilon$. Then under the isomorphisms (77), (79) the real structure on $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ given by complex conjugation corresponds to $\nu\otimes\psi^{\ast}\mapsto\bar{\nu}\otimes\bar{\psi}^{\ast}$. Now the map $s\mapsto\bar{s}$ conjugates the Killing constant of any Killing spinor and conjugates the eigenvalue of any eigenspinor of $D^{A}-\operatorname{ad}(\Phi)$. So $s\mapsto\bar{s}$ gives antilinear maps $E_{x}^{\pm}\to E_{x}^{\mp}$ and $K^{\pm}\to K^{\mp}$. Therefore the two isomorphisms (113) are exchanged by the real structure on $T^{\mathbb{C}}_{x}M^{f}_{n,p}$.
Given a non-zero Killing spinor $\psi\in K^{-}$, we define an almost complex structure on $M^{f}_{n,p}$ by declaring that $T^{1,0}_{x}M^{f}_{n,p}=E_{x}^{+}\otimes\langle\bar{\psi}\rangle$ and $T^{0,1}_{x}M^{f}_{n,p}=E_{x}^{-}\otimes\langle{\psi}\rangle$, in which $\langle\psi\rangle\subset K^{-}$ denote the span of $\psi$. The associated linear map $J_{\langle\psi\rangle}:T_{x}M^{f}_{n,p}\to T_{x}M^{f}_{n,p}$ is then given by
$J_{\langle\psi\rangle}(\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi})=-i\nu\otimes\psi+i\bar{\nu}\otimes\bar{\psi}\quad\forall\nu\in E_{x}^{-}.$ (121)
@@
To ensure that this is well-defined we must check that every real tangent vector can be written in the form $\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$, or in other words, that $\nu\mapsto\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$ defines an isomorphism $E^{-}_{x}\to T_{x}M^{f}_{n,p}$. This map is injective because if $\nu$ is in its kernel then $0=\nu\otimes\psi^{\ast}(\bar{\psi})+\bar{\nu}\otimes\bar{\psi}^{\ast}(\bar{\psi})=-\langle\psi,\psi\rangle\nu$, and $\langle\psi,\psi\rangle\neq 0$ because $\psi\neq 0$. It is an isomorphism because its domain and target have the same real dimension.
These almost complex structures $J_{\langle\psi\rangle}$ coincide with the pluricomplex structures studied by Bielawski–Schwachhofer [5] and are therefore integrable. To see this, we write them out in local coordinates. We once again work in the upper half plane model and choose $\psi=\psi_{1}$ given in (109). Then a tangent vector $(a,\phi)$ can be written
$\operatorname{cl}_{S}(a,\phi)$ $=\begin{pmatrix}\phi-i\rho a_{\rho}&-i\rho(a_{x}-ia_{y})\\ -i\rho(a_{x}+ia_{y})&\phi+i\rho a_{\rho}\end{pmatrix}=\nu\otimes\psi^{\ast}+\bar{\nu}\otimes\bar{\psi}^{\ast},$ (122)
$\nu$ $:=-i\rho^{\frac{3}{2}}\begin{pmatrix}a_{x}-ia_{y}\\ -a_{\rho}+\frac{i}{\rho}\phi\end{pmatrix}.$
The almost complex structure (121) is therefore given by
$J_{\langle\psi\rangle}:(a_{x},a_{y},a_{\rho},\phi)\mapsto(-a_{y},a_{x},-\tfrac{1}{\rho}\phi,\rho a_{\rho}).$ (123)
In terms of the complex coordinates $z=x+iy$ and $v=\rho e^{i\theta}$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$, the components of $\nu$ are proportional to the holomorphic components $b_{z},b_{v}$ of the 1-form $b=a+\phi\,d\theta$. So $J_{\langle\psi\rangle}$ is the natural almost complex structure on the moduli space of circle-invariant instantons on $\mathbb{C}^{2}$, which is the same as the complex structure studied in [5].
It is worth pointing out that, although $J_{\langle\psi\rangle}$ was defined using the gauge fixing condition (25), the real tangent vector $a+\phi=\nu\otimes\psi^{\ast}+\bar{\nu}\otimes\bar{\psi}^{\ast}$ in (121) does not satisfy (25). This is because the two summands separately satisfy the two different gauge-fixing conditions in (25), and their sum satisfies neither. The most natural way to write the gauge-fixing condition satisfied by the real tangent vector in (121) is to regard the monopole as a circle-invariant instanton on $\mathbb{R}^{4}$. In terms of the 1-form $b=a+\phi\,d\theta$ on $\mathbb{R}^{4}$, the components $b^{1,0},b^{0,1}$ satisfy the two alternative gauge-fixing conditions (25). A calculation similar to (58) in section 3 shows that these can be written in terms of the Hodge star operator $*_{E}$ of the Euclidean metric on $\mathbb{R}^{4}$ as follows:
$d^{B}*_{E}\frac{1}{v^{2}}b^{1,0}=0=d^{B}*_{E}\frac{1}{\bar{v}^{2}}b^{0,1},$ (124)
where $v=\rho\mathrm{e}^{i\theta}$. Since $*_{E}b^{1,0}$ is a form of type $(2,1)$,
$d^{B}*_{E}\frac{1}{v^{2}}b^{1,0}=\bar{\partial}^{B}*_{E}\frac{1}{v^{2}}b^{1,0}=\frac{1}{v^{2}}\bar{\partial}^{B}*_{E}b^{1,0}=\frac{1}{v^{2}}d^{B}*_{E}b^{1,0}.$ (125)
This, together with a similar calculation for $b^{0,1}$, shows that the gauge-fixing condition (25) for $a+\phi$ is equivalent to $d^{B}*_{E}b=0$, which is the usual Coulomb gauge-fixing condition on $\mathbb{R}^{4}$. Written in terms of $(a,\phi)$ it reads $\partial^{A}_{x}a_{x}+\partial^{A}_{y}a_{y}+\partial^{A}_{\rho}a_{\rho}+a_{\rho}/\rho+[\Phi,\phi]/\rho^{2}=0$. This is clearly different from the gauge-fixing condition (25), which takes the form $\partial^{A}_{x}a_{x}+\partial^{A}_{y}a_{y}+\partial^{A}_{\rho}a_{\rho}+a_{\rho}/\rho+[\Phi,\phi]/\rho^{2}\pm 2i\phi=0$.
The complex structures $J_{\langle\psi\rangle}$ depend on a line $\langle\psi\rangle\in K^{-}$, or equivalently, a point in $\mathbb{P}K^{-}$. The space $\mathbb{P}K^{-}$ can be identified with the boundary $S^{2}_{\infty}$ in a natural way. For any non-zero Killing vector $\psi$, there is a unique point $z_{0}\in S^{2}_{\infty}$ such that $|\psi|$ decays as one approaches $z_{0}$ along any geodesic. Indeed, from the explicit solutions (109) we see that
$\bar{z}_{0}\psi_{1}+\psi_{2}=\rho^{-\frac{1}{2}}\begin{pmatrix}\bar{z}_{0}-\bar{z}\\ \rho\end{pmatrix}$ (126)
decays as $\rho^{\frac{1}{2}}$ as along geodesics approaching the point $z_{0}\in\mathbb{C}\subset S^{2}_{\infty}$, and grows as $\rho^{-\frac{1}{2}}$ along other geodesics. So there is a natural anti-holomorphic identification $S^{2}_{\infty}\cong\mathbb{P}K^{-}$, and a holomorphic identification $S^{2}_{\infty}\cong\mathbb{P}K^{+}$. Thus $\langle\psi\rangle\to J_{\langle\psi\rangle}$ is a map from $S^{2}_{\infty}$ to the space of complex structures on $T_{x}M^{f}_{n,p}$. This map is holomorphic, in the sense that $T^{1,0}_{x}M^{f}_{n,p}$ depends holomorphically on $z_{0}\in S^{2}_{\infty}$. This perspective on pluricomplex structures is the starting point of the paper [5].
@@
Now let us consider the compatibility of the complex structures $J_{\langle\psi\rangle}$ with the bilinear forms $g_{\pm}$. Under the isomorphism (77), the complex-linear inner product on $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ becomes $g(s\otimes t,s^{\prime}\otimes t^{\prime})=\frac{1}{2}(s,s^{\prime})(t,t^{\prime})$. So for any $\nu\otimes\psi\in E^{-}_{x}\otimes K^{-}$,
$g_{-}(\nu\otimes\psi,\nu\otimes\psi)=-\frac{1}{4}\int_{H^{3}}(\psi,\psi)\operatorname{Tr}\big{(}\nu,\nu\big{)}\mathrm{Vol}_{H^{3}}=0.$ (127)
So $g_{-}$ is compatible with the complex structures in the sense that it makes antiholomorphic tangent vectors null. Similarly, $g_{+}$ makes holomorphic tangent vectors null. If $g_{+}$ and $g_{-}$ were real this would imply that they are compatible with $J_{\langle\psi\rangle}$ in the usual sense that $g_{\pm}(J_{\langle\psi\rangle}X,J_{\langle\psi\rangle}X)=g_{\pm}(X,X)$ for all $X$.
We have seen that the tensor product decomposition (113) induces complex structures on the moduli space. A second interesting structure is a natural decomposition of the bilinear forms $g_{\pm}$. The Killing spinor equation implies that $(\psi_{1},\psi_{2})$ is a constant function on $H^{3}$ for any $\psi_{1},\psi_{2}\in K^{\pm}$. Therefore $g_{\pm}$ can be written
$g_{\pm}(\nu_{1}\otimes\psi_{1},\nu_{2}\otimes\psi_{2})=\frac{1}{2}(\psi_{1},\psi_{2})\omega_{\pm}(\nu_{1},\nu_{2}),\quad\forall\psi_{\alpha}\in K^{\pm},\nu_{\alpha}\in E_{x}^{\pm},$ (128)
in which
$\omega_{\pm}(\nu_{1},\nu_{2}):=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}\big{(}\nu_{1},\nu_{2}\big{)}\mathrm{Vol}_{H^{3}}.$ (129)
In other words, the $L^{2}$ symmetric form $g_{\pm}$ on $T_{x}M^{f}_{n,p}$ can be decomposed into an $L^{2}$ skew-symmetric form $\omega_{\pm}$ on $E_{x}^{\pm}$ and the pointwise symplectic pairing (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $K^{\pm}$. If $g_{\pm}$ is non-degenerate, then $\omega_{\pm}$ is also nondegenerate and is a symplectic form. So if $g_{\pm}$ is non-degenerate, the structure group $SO(4n)$ of the tangent bundle can be reduced to $Sp(2n)$. This $Sp(2n)$-structure is not expected to be integrable, and it would be interesting to determine its torsion classes.
@@
From $\omega_{\pm}$, one can construct a family of holomorphic (symplectic) forms on the tangent bundle. These are parametrised by $\psi\in K^{\pm}$ and defined by
$\Omega_{\psi}^{\pm}((a_{1},\phi_{1}),(a_{2},\phi_{2})):=\omega_{\pm}((a_{1}+\phi_{1})\cdot\psi,(a_{2}+\phi_{2})\cdot\psi).$ (130)
@@
The two-forms $\omega_{\pm}$ are reminiscent of a construction due to Nash [24]. To explain this, we first recall relevant part of [24]. Nash considered the minitwistor space, which is the space of oriented geodesics in $H^{3}$ and can be be identified with $S^{2}_{\infty}\times S^{2}_{\infty}$. From our perspective it is natural to identify the minitwistor space with $\mathbb{P}K^{+}\times\mathbb{P}K^{-}$. A monopole $x\in M^{f}_{n,p}$ is identified with a spectral curve $S$ in a bundle over minitwistor space, and the tangent space $T_{x}^{\mathbb{C}}M^{f}_{n,p}$ is identified with the space of sections of the normal bundle $\hat{N}$ of this curve. Nash obtains in eq. (28) of [24] two decompositions
$T_{x}^{\mathbb{C}}M^{f}_{n,p}$ $\cong H^{0}(S,\hat{N}(-1,0))\otimes H^{0}(S,\mathcal{O}(1,0))$ (131)
$T_{x}^{\mathbb{C}}M^{f}_{n,p}$ $\cong H^{0}(S,\hat{N}(0,-1))\otimes H^{0}(S,\mathcal{O}(0,1)).$ (132)
Here $\mathcal{O}(1,0)$ and $\mathcal{O}(0,1)$ are the duals of the tautological bundles of $\mathbb{P}K^{\pm}$; the spaces $H^{0}(S,\mathcal{O}(1,0))$, $H^{0}(S,\mathcal{O}(0,1))$ of sections are naturally identified with $K^{\pm}$. So Nash’s decompositions are very similar to (113). In Lemma 3.5 Nash constructs skew-symmetric pairings on his spaces $H^{0}(S,\hat{N}(-1,0))$, $H^{0}(S,\hat{N}(0,-1))$ and shows that they are non-degenerate. We conjecture that his decomposition coincides with (113), and that these symplectic forms coincide with (129). If true, this would imply that $\omega_{\pm}$ are nondegenerate and hence, via (128), that $g_{\pm}$ are nondegenerate.
@@ 5 Calculating the metric
In this section we obtain some explicit solutions of the gauge-fixing condition (25) and hence evaluate some components of the bilinear form $g_{\pm}$. In particular, we calculate $g_{\pm}$ for the 1-monopole moduli space and confirm that it is a metric.
@@
$[X_{i},X_{j}]=-\epsilon_{ijk}X_{k},\quad[X_{i},Y_{j}]=-\epsilon_{ijk}Y_{k},\quad[Y_{i},Y_{j}]=\epsilon_{ijk}X_{k}.$ (137)
@@
$\begin{split}\nu_{1}&=-i\rho^{3/2}\begin{pmatrix}F_{xy}\\ F_{y\rho}+iF_{\rho x}\end{pmatrix},\\ \nu_{2}&=i\rho^{3/2}\begin{pmatrix}(x-iy)F_{xy}-\rho(F_{y\rho}-iF_{\rho x})\\ (x-iy)(F_{y\rho}+iF_{\rho x})+\rho F_{xy}\end{pmatrix}.\end{split}$ (145)
@@
$\begin{split}\omega_{-}(\nu_{1},\nu_{2})&=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\operatorname{Vol}_{H^{3}}\\ &=-\frac{1}{2}\int_{H^{3}}\rho^{4}\operatorname{Tr}(F_{xy}^{2}+F_{y\rho}^{2}+F_{\rho x}^{2})\operatorname{Vol}_{H^{3}}\\ &=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge*F^{A})=2\pi np,\end{split}$ (146)
@@
Equation (148) fully determines $g_{-}$ in the case of charge 1. It is well-known that the Bogmolny equations admit a spherically-symmetric solution when $n=1$. The moduli space $M^{f}_{1,p}$ can be generated by transforming this monopole using isometries of $H^{3}$ and unframed gauge transformations $g=\exp(s\Phi)$. It is therefore diffeomorphic to ×$\mathbb{R}\times SO(3,1)/SO(3)=\mathbb{R}\times H^{3}$. The bilinear form $g_{-}$ is clearly invariant under the transitive action of $SO(3,1)\times\mathbb{R}$, so it is determined by its value at a single point in the moduli space.
@@
$-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\operatorname{Vol}_{H^{3}}=\\ \int_{H^{3}}\frac{12\lambda^{4}\rho^{4}}{((x-x_{0})^{2}+(y-y_{0})^{2}+\rho^{2}+\lambda^{2})^{4}}\frac{dx\,dy\,d\rho}{\rho^{3}}=\pi,$ (160)
@@
$\chi_{1}=i\frac{\rho\sigma_{1}-x\sigma_{3}}{1+|\mathbf{x}|^{2}}.$ (173)
@@
$\chi_{2}=i\frac{\rho\sigma_{2}-y\sigma_{3}}{1+|\mathbf{x}|^{2}},\quad\chi_{3}=\frac{i\sigma_{3}}{1+|\mathbf{x}|^{2}}+\Phi,$ (176)

### v2

@@ 4.3 Symplectic and pluricomplex structures
We extend complex conjugation to $\mathfrak{su}(2)$-valued spinors trivially, that is, for $s\in S$, $\upsilon\in\mathfrak{su}(2)$ we define $\overline{s\otimes\upsilon}=\overline{s}\otimes\upsilon$. Then under the isomorphisms (110), (112) the real structure on $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ given by complex conjugation corresponds to $\nu\otimes\psi^{\ast}\mapsto\bar{\nu}\otimes\bar{\psi}^{\ast}$.
Now the map $s\mapsto\bar{s}$ conjugates the Killing constant of any Killing spinor and conjugates the eigenvalue of any eigenspinor of $D^{A}-\operatorname{ad}(\Phi)$. So $s\mapsto\bar{s}$ gives antilinear maps $E_{x}^{\pm}\to E_{x}^{\mp}$ and $K^{\pm}\to K^{\mp}$. Therefore the two isomorphisms (146) are exchanged by the real structure on $T^{\mathbb{C}}_{x}M^{f}_{n,p}$.
Proof of Proposition 29.
From the discussion above, $\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$ is a real tangent vector for any $\nu\in E_{x}^{+}$. The linear map $J_{\langle\psi\rangle}:T_{x}M^{f}_{n,p}\to T_{x}M^{f}_{n,p}$ will be given by
$J_{\langle\psi\rangle}(\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi})=i\nu\otimes\psi-i\bar{\nu}\otimes\bar{\psi}\quad\forall\nu\in E_{x}^{+}.$ (157)
@@
To ensure that this is well-defined we must check that every real tangent vector can be written in the form $\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$, or in other words, that $\nu\mapsto\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$ defines an isomorphism $E^{-}_{x}\to T_{x}M^{f}_{n,p}$. This map is injective because if $\nu$ is in its kernel then $0=\nu\otimes\psi^{\ast}(\bar{\psi})+\bar{\nu}\otimes\bar{\psi}^{\ast}(\bar{\psi})=-\langle\psi,\psi\rangle\nu$, and $\langle\psi,\psi\rangle\neq 0$ because $\psi\neq 0$. It is an isomorphism because its domain and target both have real dimension $4n$. ∎
Recall that an almost complex structure $J$ is compatible with a riemannian metric $g$ if and only if every holomorphic tangent vector $X$ is null, i.e. $g(X,X)=0$. The almost complex structures $J_{\langle\psi\rangle}$ defined above are compatible with the bilinear pairing $g_{+}$ in a similar way:
Proposition 30.
Let $x\in M^{f}_{n,p}$ and let $\langle\psi\rangle\in K^{\pm}$. Then for all $X\in T_{x}M^{f}_{n,p}$,
$J_{\langle\psi\rangle}X=iX\implies g_{+}(X,X)=0.$ (158)
Proof.
@@
By definition, $J_{\langle\psi\rangle}X=iX$ if and only if $X=\nu\otimes\psi$ for some $\nu\in E^{+}_{x}$. Under the isomorphism (110), the complex-linear inner product on $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ becomes $g(s\otimes t,s^{\prime}\otimes t^{\prime})=\frac{1}{2}(s,s^{\prime})(t,t^{\prime})$. So for any $\nu\otimes\psi\in E^{+}_{x}\otimes K^{+}$,
$g_{+}(\nu\otimes\psi,\nu\otimes\psi)=-\frac{1}{4}\int_{H^{3}}(\psi,\psi)\operatorname{Tr}\big{(}\nu,\nu\big{)}\mathrm{Vol}_{H^{3}}=0.$ (159)
We have seen that the tensor product decomposition (146) induces almost complex structures on the moduli space. The decomposition also gives rise to tensor product decompositions of $g_{\pm}$, as we now explain.
Recall that $(\cdot,\cdot)$ denotes the symplectic form on the spinor bundle of $H^{3}$. From this, we define $L^{2}$ skew bilinear forms $\omega_{E^{\pm}}$ on the spaces $E_{x}^{\pm}$ of eigenspinors:
$\omega_{E^{\pm}}(\nu_{1},\nu_{2}):=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}\big{(}\nu_{1},\nu_{2}\big{)}\mathrm{Vol}_{H^{3}}.$ (160)
For the Killing spinors $K_{\pm}$, we have the following result:
Lemma 31.
Let $\psi_{1},\psi_{2}$ be two Killing spinors on $H^{3}$ with the same Killing constant. Then $(\psi_{1},\psi_{2})$ is a constant function.
Proof.
By the Killing spinor equation, for any tangent vector $X$:
$\nabla_{X}(\psi_{1},\psi_{2})$ $=(\nabla_{X}\psi_{1},\psi_{2})+(\psi_{1},\nabla_{X}\psi_{2})$ (161)
$=\pm\frac{i}{2}\big{[}(X\cdot\psi_{1},\psi_{2})+(\psi_{1},X\cdot\psi_{2})\big{]}=0.$ (162)
For any $\psi_{1},\psi_{2}\in K^{\pm}$ we define $\omega_{K^{\pm}}(\psi_{1},\psi_{2})\in\mathbb{C}$ to be the value of the constant function $(\psi_{1},\psi_{2})$ on $H^{3}$. Then $\omega_{K^{\pm}}$ is a symplectic form on $K^{\pm}$. We then have:
Proposition 32.
The pairings $g_{\pm},\omega_{E^{\pm}},\omega_{K^{\pm}}$ on $T^{\mathbb{C}}M^{f}_{n,p}$, $E^{\pm}$ and $K^{\pm}$ are related as follows:
$g_{\pm}=\frac{1}{2}\omega_{K^{\pm}}\otimes\omega_{E^{\pm}}.$ (163)
Proof.
Under the isomorphism (110), the riemannian inner product on $TH^{3}$ corresponds to the symmetric pairing $\frac{1}{2}(\cdot,\cdot)\otimes(\cdot,\cdot)$ on $S\otimes S$. So for any $\psi_{\alpha}\in K^{\pm},\nu_{\alpha}\in E_{x}^{\pm}$,
$g_{\pm}(\nu_{1}\otimes\psi_{1},\nu_{2}\otimes\psi_{2})$ $=-\frac{1}{4}\int_{H^{3}}(\psi_{1},\psi_{2})\operatorname{Tr}(\nu_{1},\nu_{2})\,\mathrm{Vol}_{H^{3}}$ (164)
$=-\frac{1}{4}(\psi_{1},\psi_{2})\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\,\mathrm{Vol}_{H^{3}}$ (165)
$=\frac{1}{2}\omega_{K^{\pm}}(\psi_{1},\psi_{2})\omega_{E^{\pm}}(\nu_{1},\nu_{2}),$ (166)
where we used the fact that $(\psi_{1},\psi_{2})$ is constant. ∎
The skew forms $\omega_{E^{\pm}}$ are reminiscent of a construction due to Nash [26]. To explain this, we first recall relevant part of [26]. Nash worked in minitwistor space, which is the space of oriented geodesics in $H^{3}$ and can be identified with an open subset of $S^{2}_{\infty}\times S^{2}_{\infty}$. From our perspective it is natural to identify this with $\mathbb{P}K^{+}\times\mathbb{P}K^{-}$. A hyperbolic monopole corresponds to a spectral curve in minitwistor space such that a certain holomorphic bundle $L$ is trivial over this curve. Nash defines a framed moduli space $M^{t}_{n,p}$ to be the space of spectral curves $\hat{S}$ in the total space of the bundle $L$ over minitwistor space. He identifies his tangent space $T_{x}^{\mathbb{C}}M^{t}_{n,p}$ with the space of sections of the normal bundle $\hat{N}$ of this curve. Nash obtains in eq. (28) of [26] two decompositions
$T_{x}^{\mathbb{C}}M^{t}_{n,p}$ $\cong H^{0}(S,\hat{N}(-1,0))\otimes H^{0}(S,\mathcal{O}(1,0))$ (167)
$T_{x}^{\mathbb{C}}M^{t}_{n,p}$ $\cong H^{0}(S,\hat{N}(0,-1))\otimes H^{0}(S,\mathcal{O}(0,1)).$
Here $\mathcal{O}(1,0)$ and $\mathcal{O}(0,1)$ are the duals of the tautological bundles of $\mathbb{P}K^{\pm}$; the spaces $H^{0}(S,\mathcal{O}(1,0))$, $H^{0}(S,\mathcal{O}(0,1))$ of sections are naturally identified with $K^{\pm}$. So Nash’s decompositions are very similar to (146). In Lemma 3.5 Nash constructs skew-symmetric pairings on his spaces $H^{0}(S,\hat{N}(-1,0))$, $H^{0}(S,\hat{N}(0,-1))$ and shows that they are in fact symplectic forms. We conjecture that his framed moduli space $M^{t}_{n,p}$ is naturally locally diffeomorphic to our $M^{f}_{n,p}$, that his decomposition (167) coincides with (146), and that his symplectic forms coincide with (160). If true, this would imply that $\omega_{E^{\pm}}$ are nondegenerate and hence, via (163), that $g_{\pm}$ are nondegenerate. It would also give a reduction of the structure group of $T^{\mathbb{C}}M^{f}_{n,p}$ to $Sp(2n,\mathbb{C})$.
Remark 33.
@@
From $\omega_{E^{\pm}}$, one can construct a family of holomorphic (symplectic) forms on the tangent bundle. These are parametrised by $\psi\in K^{\pm}$ and defined by
$\Omega_{\psi}^{\pm}((a_{1},\phi_{1}),(a_{2},\phi_{2})):=\omega_{E^{\pm}}((a_{1}+\phi_{1})\cdot\psi,(a_{2}+\phi_{2})\cdot\psi).$ (168)
@@ 5 Calculating the metric
In this section we obtain some explicit solutions of the gauge-fixing condition (49) and hence evaluate some components of the bilinear form $g_{\pm}$. In particular, we calculate $g_{\pm}$ for the 1-monopole moduli space and confirm that it is a riemannian metric.
@@
$X_{i},X_{j}]=-\epsilon_{ijk}X_{k},\quad[X_{i},Y_{j}]=-\epsilon_{ijk}Y_{k},\quad[Y_{i},Y_{j}]=\epsilon_{ijk}X_{k}.$ (173)
@@
[…]
@@
$\begin{split}\omega_{E^{-}}(\nu_{1},\nu_{2})&=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\operatorname{Vol}_{H^{3}}\\ &=-\frac{1}{2}\int_{H^{3}}\rho^{4}\operatorname{Tr}(F_{xy}^{2}+F_{y\rho}^{2}+F_{\rho x}^{2})\operatorname{Vol}_{H^{3}}\\ &=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge*F^{A})=2\pi np,\end{split}$ (182)
@@
Equation (184) fully determines $g_{-}$ in the case of charge 1. It is well-known that the Bogmolny equations admit a spherically-symmetric solution when $n=1$. The moduli space $M^{f}_{1,p}$ can be generated by transforming this monopole using isometries of $H^{3}$ and unframed gauge transformations $g=\exp(s\Phi)$. It is therefore diffeomorphic to $\mathbb{R}\times SO(3,1)/SO(3)=\mathbb{R}\times H^{3}$. The bilinear form $g_{-}$ is clearly invariant under the transitive action of $SO(3,1)\times\mathbb{R}$, so it is determined by its value at a single point in the moduli space.
@@
$\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\operatorname{Vol}_{H^{3}}=\\ \int_{H^{3}}\frac{12\lambda^{4}\rho^{4}}{((x-x_{0})^{2}+(y-y_{0})^{2}+\rho^{2}+\lambda^{2})^{4}}\frac{dx\,dy\,d\rho}{\rho^{3}}=\pi,$ (196)
@@
$\chi_{1}=\frac{i(\rho\sigma_{1}-x\sigma_{3})}{1+|\mathbf{x}|^{2}}.$ (209)
@@
$\chi_{2}=\frac{i(\rho\sigma_{2}-y\sigma_{3})}{1+|\mathbf{x}|^{2}},\quad\chi_{3}=\frac{i\sigma_{3}}{1+|\mathbf{x}|^{2}}+\Phi,$ (212)